OpenStudy (anonymous):

slope of the tangent to the curve at y-cos2x+1 at x=pi/6 --- PLEASE HELP!!!!

OpenStudy (dan815):

y=cos2x+1 at x=pi/6

OpenStudy (dan815):

like that?

OpenStudy (dan815):

y'=2sin2x sub x=pi/6 and solve

OpenStudy (anonymous):

i tried, i dont know how.